# Beyond Local Sharpness: Communication-Efficient Global Sharpness-aware Minimization

### v1

@@ Abstract
Federated learning (FL) enables collaborative model training with privacy preservation. Data heterogeneity across edge devices (clients) can cause models to converge to sharp minima, negatively impacting generalization and robustness. Recent approaches use client-side sharpness-aware minimization (SAM) to encourage flatter minima, but the discrepancy between local and global loss landscapes often undermines their effectiveness, as optimizing for local sharpness does not ensure global flatness. This work introduces FedGloSS (Federated Global Server-side Sharpness), a novel FL approach that prioritizes the optimization of global sharpness on the server, using SAM. To reduce communication overhead, FedGloSS cleverly approximates sharpness using the previous global gradient, eliminating the need for additional client communication. Our extensive evaluations demonstrate that FedGloSS consistently reaches flatter minima and better performance compared to state-of-the-art FL methods across various federated vision benchmarks.
@@ 1 Introduction
Federated Learning (FL) (McMahan et al., 2017) provides a powerful framework to collaboratively train machine learning models on private data distributed across multiple endpoints. Unlike traditional methods, FL enables edge devices (clients), like smartphones or IoT (Internet of Things) hardware, to train a shared model without compromising their sensitive information. This is achieved through communication rounds, where clients independently train on their local data and then exchange updated model parameters with a central server, preserving data privacy. The optimization on the server side relies on pseudo-gradients (Reddi et al., 2021), i.e., the average of the difference between the global model and the client’s update, which serve as an estimate of the true global gradient on the overall dataset. This approach holds immense potential for privacy-sensitive applications, proving its value in areas like healthcare (Liu et al., 2021; Antunes et al., 2022; Rauniyar et al., 2023; Nevrataki et al., 2023), finance (Nevrataki et al., 2023), autonomous driving (Fantauzzo et al., 2022; Shenaj et al., 2023; Miao et al., 2023), IoT (Zhang et al., 2022), and more (Li et al., 2020a; Wen et al., 2023). However, the real-world deployment of FL presents unique challenges stemming from data heterogeneity and communication costs (Li et al., 2020b). Clients gather their data influenced by various factors such as personal habits or geographical locations, leading to inherent differences across devices (Kairouz et al., 2021; Hsu et al., 2020; Shenaj et al., 2023). This results in the global model suffering from degraded performance and slower convergence (Li et al., 2020d; Karimireddy et al., 2020a; b; Caldarola et al., 2022), with instability emerging as client-specific optimization paths diverge from the global one. This phenomenon, known as client drift (Karimireddy et al., 2020b), limits the model’s ability to generalize to the overall underlying distribution.
@@
While many FL approaches focus on mitigating client drift through client-side regularization (Li et al., 2020c; Acar et al., 2021; Varno et al., 2022), a recent trend leverages the geometry of the loss landscape to improve generalization (Caldarola et al., 2022; Qu et al., 2022; Sun et al., 2023b; Dai et al., 2023; Sun et al., 2023a). These methods build upon the notion that convergence to sharp minima correlates with poor generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Jiang et al., 2019). FedSam (Caldarola et al., 2022; Qu et al., 2022) employs Sharpness-aware Minimization (Sam) (Foret et al., 2021) in local training to guide clients toward flatter loss regions, enhancing global performance. This comes at the cost of increased client-side computation, since Sam requires two forward/backward passes for each local optimization step: a gradient ascent step to compute the maximum sharpness and a descent step for sharpness and loss value minimization. Although FedSam and its variants (Sun et al., 2023b; Dai et al., 2023) demonstrated their effectiveness in various settings, they rely solely on local flatness, assuming that minimizing sharpness locally leads to a globally flat minimum. However, in real-world scenarios with significant data heterogeneity, there can be substantial discrepancies between local and global loss landscapes. As a consequence, optimizing for local sharpness does not guarantee the global model will reside in a flat region (Fig. 1). Addressing these limitations, FedSmoo (Sun et al., 2023a) uses the alternating direction method of multipliers (ADMM) (Boyd et al., 2011) to include global sharpness information in Sam’s local training. While this approach reduces the inconsistency between local and global geometries, it increases communication cost by requiring double the bandwidth in each round. This hinders its real-world applicability, as FL relies on minimizing communication overhead (i.e., both message size and exchange frequency) to avoid network congestion and account for potential connection failures, that are common in practical deployments.
Given the limitations of existing methods, achieving convergence to global flat minima while maintaining communication efficiency in heterogeneous FL remains a critical challenge. To address this, we propose FedGloSS (Federated Global Server-side Sharpness) that directly optimizes global sharpness by using Sam on the server side, avoiding additional exchanges over the network. Such adaptation is not straightforward, as Sam would require dual exchanges with each client set per round to solve its optimization problem. Instead, FedGloSS approximates the sharpness measure using available previous pseudo-gradients. As a result, FedGloSS facilitates faster training and keeps communication efficiency. To summarize, our core contributions are the following:
@@
  • We show the importance of aligning global and local solutions and illustrate how Sam, especially on the server side, enables effective ADMM use in FL. While typically ADMM-based methods suffer from parameter explosion Varno et al. (2022), we show that by targeting flat minima, Sam encourages smaller gradient steps and minimal weight updates, leading to a significantly more stable algorithm.
@@ 2 Related works
In the last few years, Federated Learning (FL) (McMahan et al., 2017) garnered significant attention from both the machine learning and computer vision communities. While the former has primarily focused on optimizing FL algorithms and guaranteeing their convergence (Li et al., 2020d; Acar et al., 2021; Reddi et al., 2021), the latter has explored its applications in real-world settings, spanning diverse domains like autonomous driving (Fantauzzo et al., 2022; Shenaj et al., 2023; Miao et al., 2023) and healthcare (Liu et al., 2021). The key appeal of FL lies in its ability to efficiently learn from privacy-protected, distributed data while complying with regulations and leveraging edge resources. Real-world deployments of FL range across both cross-silo and cross-device settings (Kairouz et al., 2021). This work focuses on the latter, with up to millions of individual devices at the network edge, with typically limited data and computational power, and potential unavailability due to battery life or network connectivity issues. User-specific factors like geographical location, capturing devices and daily habits introduce inherent bias and statistical heterogeneity into the local datasets. In this setting, FedGloSS aims to learn a global model that generalizes to the overall data distribution under statistical heterogeneity without increasing communication complexity, unlike other algorithms for local-global consistency in heterogeneous FL.
@@
Flatness search in FL. Recent research has explored the connection between loss landscape geometry and generalization in heterogeneous FL. Studies suggest that convergence to sharp minima might hinder generalization performance (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Petzka et al., 2021). Sam (Sharpness-Aware Minimization) (Foret et al., 2021) tackles this issue by guiding the optimization toward flatter regions, seeking minima that exhibit both low loss and low sharpness. FedSam (Caldarola et al., 2022; Qu et al., 2022) deploys Sam in local training, marking the first step toward leveraging loss surface geometry in FL to reduce discrepancies between local and global objectives, ultimately improving the global model’s generalization ability. Following its success, FedSpeed (Sun et al., 2023b) uses perturbed gradients as Sam to reduce local overfitting, FedGamma (Dai et al., 2023) combines the stochastic variance reduction of Scaffold with Sam and Shi et al. (2023) show FedSam’s effectiveness in mitigating the negative effects of differential privacy. However, these approaches rely on local sharpness information, assuming its minimization directly translates to a globally flat minimum. This may not always be true, as we hypothesize discrepancies may exist between the geometries of local and global losses. Optimizing local sharpness alone does not guarantee a server model residing in a flat region of the global loss landscape (Fig. 1). Addressing these limitations, FedSmoo (Sun et al., 2023a) applies ADMM (Boyd et al., 2011) to the sharpness measure to enforce global and local consistency. This adds communication overhead, doubling the message size in each round and hindering its real-world practicality. In contrast, our work focuses on minimizing global sharpness while maintaining communication efficiency. Lastly, building on Stochastic Weight Averaging (Izmailov et al., 2018), other works (Caldarola et al., 2022; 2023) use a window-based average of global models across rounds to reach wider minima. Being agnostic to the underlying optimization algorithm, they remain orthogonal to our approach.
@@
Heterogeneity in FL. The de-facto standard algorithm for FL is FedAvg (McMahan et al., 2017), which updates the global model with a weighted average of the clients’ parameters. However, FedAvg struggles when faced with heterogeneous data distributions, leading to performance degradation and slow convergence due to the local optimization paths diverging from the global one (Karimireddy et al., 2020b). Reddi et al. (2021) shows FedAvg is equivalent to applying Stochastic Gradient Descent (Sgd) (Ruder, 2016) with a unitary learning rate on the server side, using the difference between the initial global model parameters and the clients’ updates as pseudo-gradient, opening the door to alternative optimizers beyond Sgd to improve performance and convergence speed. Building on this intuition, this work proposes Sam (Foret et al., 2021) as a server-side optimizer to enhance generalization by converging toward global flat minima. Since Sam requires two optimization steps per iteration, a direct adaptation to the FL setting would double communication exchanges between clients and server; FedGloSS overcomes this limitation and maintains communication efficiency through the use of the latest pseudo-gradient as sharpness approximation.
@@
Several approaches address client drift by adding regularization during local training. FedProx (Li et al., 2020c) introduces a term to keep local parameters close to the global model, FedDyn (Acar et al., 2021) employs ADMM to align local and global convergence points, AdaBest (Varno et al., 2022) adjusts local updates with an adaptive bias estimate, and Scaffold (Karimireddy et al., 2020b) applies stochastic variance reduction. Momentum-based techniques (Sutskever et al., 2013) are also employed to maintain a consistent global trajectory, either on the server side (e.g., FedAvgM (Hsu et al., 2019)) or by incorporating global information into local training (Karimireddy et al., 2020a; Kim et al., 2022; Gao et al., 2022; Zaccone et al., 2023). Unlike FedDyn, where ADMM can lead to parameter explosion (Varno et al., 2022), our FedGloSS successfully leverages ADMM to align global and local solutions, even under extreme heterogeneity, aided by Sam on server.
@@
Centralized SAM. To avoid doubling client-server exchanges caused by Sam’s two-step process, FedGloSS draws on insights from the literature on Sam in centralized settings. Several strategies have been proposed to minimize computational overhead, including reducing the number of parameters needed to compute the sharpness-aware components (Du et al., 2022a), or approximating them (Liu et al., 2022; Du et al., 2022b; Park et al., 2023). DP-SAT (Park et al., 2023) approximates the ascent step with the gradient from the previous iteration, and SAF (Du et al., 2022b) replaces Sam’s sharpness approximation with the trajectory of weights learned during training. Aiming to the same goal, FedGloSS approximates the sharpness measure with the pseudo-gradient from the previous round on the server side, without incurring in unnecessary exchanges with the clients and effectively guiding the optimization toward globally flat minima.
@@ 3 Background
This section introduces the FL problem setting and preliminary notations on Sam (Foret et al., 2021) and FedSam (Caldarola et al., 2022; Qu et al., 2022).
@@ 3.1 Problem setting
$\min_{\operatorname{\boldsymbol{w}}}\left\{f(\operatorname{\boldsymbol{w}})=\frac{1}{C}\sum_{k\in\operatorname{\mathcal{C}}}f_{k}(\operatorname{\boldsymbol{w}})\right\},f_{k}(\operatorname{\boldsymbol{w}})\triangleq\mathbb{E}f_{k}(\operatorname{\boldsymbol{w}},\xi_{k}),$ (1)
where $C\triangleq|\operatorname{\mathcal{C}}|$ is the total number of clients, $f_{k}$ is the empirical loss on the $k$-th client (e.g., cross-entropy loss) and $\xi_{k}$ is the data sample randomly drawn from the local data distribution $\mathit{D}_{k}$. The training process is a two-phase optimization approach within each round $t\in[T]$. First, due to potential client unavailability, a subset of selected clients $\operatorname{\mathcal{C}}^{t}\subset\operatorname{\mathcal{C}}$ trains the received global model using their local optimizer ClientOpt (e.g., Sgd, Sam). Then, the server aggregates their updates with a server optimizer, ServerOpt. FedOpt (Reddi et al., 2021) solves Eq. 1 as
@@
$\operatorname{\boldsymbol{w}}^{t+1}$ $\leftarrow\operatorname{\boldsymbol{w}}^{t}-\textsc{ServerOpt}(\operatorname{\boldsymbol{w}}^{t},\Delta^{t}_{\operatorname{\boldsymbol{w}}},\eta_{s}),$ (3)
where $\Delta^{t}_{\operatorname{\boldsymbol{w}}}$ is the global pseudo-gradient at round $t$, $N=\sum_{k\in\operatorname{\mathcal{C}}^{t}}N_{k}$ the total number of images seen during the current round, $\eta_{s}$ the server learning rate, $\operatorname{\boldsymbol{w}}^{t}$ the global model and $\operatorname{\boldsymbol{w}}^{t}_{k}$ the local update resulting from training on client $k$’s data with ClientOpt for $E$ epochs. FedAvg (McMahan et al., 2017) computes $\operatorname{\boldsymbol{w}}^{t+1}$ as $\sum_{k\in\operatorname{\mathcal{C}}^{t}}\nicefrac{{N_{k}}}{{N}}\operatorname{\boldsymbol{w}}^{t}_{k}$, corresponding to one SGD step on the pseudo-gradient $\Delta^{t}_{\operatorname{\boldsymbol{w}}}$ with $\eta_{s}=1$ (Reddi et al., 2021).
@@ 3.2 Sharpness-aware Minimization
Sam (Foret et al., 2021) jointly minimizes the loss value and the sharpness of the loss landscape by solving the min-max problem
@@
$\operatorname*{arg\,max}_{\|\operatorname{\boldsymbol{\epsilon}}\|\leq\rho}f(\operatorname{\boldsymbol{w}})+\operatorname{\boldsymbol{\epsilon}}^{\top}\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})=\rho\frac{\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})}{\|\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})\|}\triangleq\operatorname{\hat{\boldsymbol{\epsilon}}}(\operatorname{\boldsymbol{w}}).$ (5)

### v2

@@ Abstract
Federated learning (FL) enables collaborative model training with privacy preservation. Data heterogeneity across edge devices (clients) can cause models to converge to sharp minima, negatively impacting generalization and robustness. Recent approaches use client-side sharpness-aware minimization (SAM) to encourage flatter minima, but the discrepancy between local and global loss landscapes often undermines their effectiveness, as optimizing for local sharpness does not ensure global flatness. This work introduces FedGloSS (Federated Global Server-side Sharpness), a novel FL approach that prioritizes the optimization of global sharpness on the server, using SAM. To reduce communication overhead, FedGloSS cleverly approximates sharpness using the previous global gradient, eliminating the need for additional client communication. Our extensive evaluations demonstrate that FedGloSS consistently achieves flatter minima and better performance compared to state-of-the-art FL methods in various federated vision benchmarks. Code available at github.com/pietrocagnasso/fedgloss.
@@ 1 Introduction
Federated Learning (FL) [39] provides a powerful framework to collaboratively train machine learning models on private data distributed across multiple endpoints. Unlike traditional methods, FL enables edge devices (clients), like smartphones or IoT (Internet of Things) hardware, to train a shared model without compromising their sensitive information. This is achieved through communication rounds, where clients independently train on their local data and then exchange updated model parameters with a central server, preserving data privacy. The optimization on the server side relies on pseudo-gradients [47], i.e., the average of the difference between the global model and the client’s update, which serve as an estimate of the true global gradient on the overall dataset. This approach holds immense potential for privacy-sensitive applications, proving its value in areas like healthcare [37, 2, 46, 41], finance [41], autonomous driving [11, 51, 40], IoT [62], and more [32, 57]. However, the real-world deployment of FL presents unique challenges stemming from data heterogeneity and communication costs [34]. Clients gather their data influenced by various factors such as personal habits or geographical locations, leading to inherent differences across devices [24, 20, 51]. This results in the global model suffering from degraded performance and slower convergence [36, 25, 26, 4], with instability emerging as client-specific optimization paths diverge from the global one. This phenomenon, known as client drift [26], limits the model’s ability to generalize to the overall underlying distribution.
@@
While many FL approaches focus on mitigating client drift through client-side regularization [35, 1, 56], a recent trend leverages the geometry of the loss landscape to improve generalization [4, 45, 54, 7, 53]. These methods build upon the notion that convergence to sharp minima correlates with poor generalization [17, 27, 23]. FedSam [4, 45] employs Sharpness-aware Minimization (Sam) [12] in local training to guide clients toward flatter loss regions, enhancing global performance. This comes at the cost of increased client-side computation, since Sam requires two forward/backward passes for each local optimization step: a gradient ascent step to compute the maximum sharpness and a descent step for sharpness and loss value minimization. Although FedSam and its variants [54, 7] demonstrated their effectiveness in various settings, they rely solely on local flatness, assuming that minimizing sharpness locally leads to a globally flat minimum [^1] We use the term “global flat minima” to refer to local minima within the global (i.e., server-side) loss landscape.. However, in real-world scenarios with significant data heterogeneity, there can be substantial discrepancies between local and global loss landscapes and optimizing for local sharpness does not guarantee the global model will reside in a flat region (Fig. 1). Addressing these limitations, FedSmoo [53] uses the alternating direction method of multipliers (ADMM) [3] to include global sharpness information in Sam’s local training. While this approach reduces the inconsistency between local and global geometries, it increases communication costs by requiring double the bandwidth in each round. This hinders its real-world applicability, as FL relies on minimizing communication overhead (i.e., message size and exchange frequency) to avoid network congestion and account for potential connection failures.
Given the limitations of existing methods, achieving global flat minima while ensuring communication efficiency in heterogeneous FL remains a critical challenge. To address this, we propose FedGloSS (Federated Global Server-side Sharpness) that directly optimizes global sharpness by using Sam on the server side, avoiding additional exchanges over the network. Such adaptation is not straightforward, as Sam would require dual exchanges with each client set per round to solve its optimization problem. Instead, FedGloSS approximates the sharpness measure using available previous pseudo-gradients. As a result, it facilitates faster training and keeps communication efficiency. To summarize, our core contributions are the following:
@@
  • We show the importance of aligning global and local solutions and illustrate how Sam, especially on the server side, enables effective ADMM use in FL. While typically ADMM-based methods suffer from parameter explosion [56], we show that by targeting flat minima, Sam encourages smaller gradient steps and minimal weight updates, leading to a significantly more stable algorithm.
@@ 2 Related works
In the last few years, Federated Learning (FL) [39] garnered significant attention from both the machine learning and computer vision communities. While the former has primarily focused on optimizing FL algorithms and guaranteeing their convergence [36, 1, 47], the latter has explored its applications in real-world settings, spanning diverse domains like autonomous driving [11, 51, 40] and healthcare [37]. The key appeal of FL lies in its ability to efficiently learn from privacy-protected, distributed data while complying with regulations and leveraging edge resources. Real-world deployments of FL range across both cross-silo and cross-device settings [24]. This work focuses on the latter, with up to millions of individual devices at the network edge, with typically limited data and computational power, and potential unavailability due to battery life or network connectivity issues. User-specific factors like geographical location, capturing devices and daily habits introduce inherent bias and statistical heterogeneity into the local datasets. In this setting, FedGloSS aims to learn a global model that generalizes to the overall data distribution under statistical heterogeneity without increasing communication complexity, unlike other algorithms for local-global consistency in heterogeneous FL.
@@
Flatness search in FL. Recent research has explored the connection between loss landscape geometry and generalization in heterogeneous FL. Studies suggest that convergence to sharp minima might hinder generalization performance [17, 27, 44]. Sam (Sharpness-Aware Minimization) [12] tackles this issue by guiding the optimization toward flatter regions, seeking minima that exhibit both low loss and low sharpness. FedSam [4, 45] deploys Sam in local training, marking the first step toward leveraging loss surface geometry in FL to reduce discrepancies between local and global objectives, ultimately improving the global model’s generalization ability. Following its success, FedSpeed [54] uses perturbed gradients as Sam to reduce local overfitting, FedGamma [7] combines the stochastic variance reduction of Scaffold with Sam and Shi et al. [52] show FedSam’s effectiveness in mitigating the negative effects of differential privacy. However, these approaches rely on local sharpness information, assuming its minimization directly translates to a globally flat minimum. This may not always be true, as we hypothesize discrepancies may exist between the geometries of local and global losses. Optimizing local sharpness alone does not guarantee a server model residing in a flat region of the global loss landscape (Fig. 1). Addressing these limitations, FedSmoo [53] applies ADMM [3] to the sharpness measure to enforce global and local consistency. This adds communication overhead, doubling the message size in each round and hindering its real-world practicality. In contrast, our work focuses on minimizing global sharpness while maintaining communication efficiency. Lastly, building on Stochastic Weight Averaging [22], other works [4, 5] use a window-based average of global models across rounds to reach wider minima. Being agnostic to the underlying optimization algorithm, they remain orthogonal to our approach.
@@
Heterogeneity in FL. The de-facto standard algorithm for FL is FedAvg [39], which updates the global model with a weighted average of the clients’ parameters. However, FedAvg struggles when faced with heterogeneous data distributions, leading to performance degradation and slow convergence due to the local optimization paths diverging from the global one [26]. Reddi et al. [47] shows FedAvg is equivalent to applying Stochastic Gradient Descent (Sgd) [49] with a unitary learning rate on the server side, using the difference between the initial global model parameters and the clients’ updates as pseudo-gradient, opening the door to alternative optimizers beyond Sgd to improve performance and convergence speed. Building on this intuition, this work proposes Sam [12] as a server-side optimizer to enhance generalization by converging toward global flat minima. Since Sam requires two optimization steps per iteration, a direct adaptation to the FL setting would double communication exchanges between clients and server; FedGloSS overcomes this limitation and maintains communication efficiency through the use of the latest pseudo-gradient as sharpness approximation.
@@
Several approaches address client drift by adding regularization during local training. FedProx [35] introduces a term to keep local parameters close to the global model, FedDyn [1] employs ADMM to align local and global convergence points, AdaBest [56] adjusts local updates with an adaptive bias estimate, and Scaffold [26] applies stochastic variance reduction. Momentum-based techniques [55] are also employed to maintain a consistent global trajectory, either on the server side (e.g., FedAvgM [19]) or by incorporating global information into local training [25, 28, 13, 61]. Unlike FedDyn, where ADMM can lead to parameter explosion [56], our FedGloSS successfully leverages ADMM to align global and local solutions, even under extreme heterogeneity, aided by Sam on server.
@@
Centralized SAM. To avoid doubling client-server exchanges caused by Sam’s two-step process, FedGloSS draws on insights from the literature on Sam in centralized settings. Several strategies have been proposed to minimize computational overhead, including reducing the number of parameters needed to compute the sharpness-aware components [9], or approximating them [38, 10, 42]. DP-SAT [42] approximates the ascent step with the gradient from the previous iteration, and SAF [10] replaces Sam’s sharpness approximation with the trajectory of weights learned during training. Aiming to the same goal, FedGloSS approximates the sharpness measure with the pseudo-gradient from the previous round on the server side, without incurring in unnecessary exchanges with the clients and effectively guiding the optimization toward globally flat minima.
@@ 3 Background
This section introduces the FL problem setting and preliminary notations on Sam [12] and FedSam [4, 45].
@@ 3.1 Problem setting
$\min_{\operatorname{\boldsymbol{w}}}\left\{f(\operatorname{\boldsymbol{w}})=\frac{1}{C}\sum_{k\in\operatorname{\mathcal{C}}}f_{k}(\operatorname{\boldsymbol{w}})\right\},f_{k}(\operatorname{\boldsymbol{w}})\triangleq\mathbb{E}f_{k}(\operatorname{\boldsymbol{w}},\xi_{k}),\vspace{-5pt}$ (1)
where $C\triangleq|\operatorname{\mathcal{C}}|$ is the total number of clients, $f_{k}$ is the empirical loss on the $k$-th client (e.g., cross-entropy loss) and $\xi_{k}$ is the data sample randomly drawn from the local data distribution $\mathit{D}_{k}$. The training process is a two-phase optimization approach within each round $t\in[T]$. First, due to potential client unavailability, a subset of selected clients $\operatorname{\mathcal{C}}^{t}\subset\operatorname{\mathcal{C}}$ trains the received global model using their local optimizer ClientOpt (e.g., Sgd, Sam). Then, the server aggregates their updates with a server optimizer, ServerOpt. FedOpt [47] solves Eq. 1 as
@@
$\operatorname{\boldsymbol{w}}^{t+1}$ $\leftarrow\operatorname{\boldsymbol{w}}^{t}-\textsc{ServerOpt}(\operatorname{\boldsymbol{w}}^{t},\Delta^{t}_{\operatorname{\boldsymbol{w}}},\eta_{s}),\vspace{-5pt}$ (3)
where $\Delta^{t}_{\operatorname{\boldsymbol{w}}}$ is the global pseudo-gradient at round $t$, $N=\sum_{k\in\operatorname{\mathcal{C}}^{t}}N_{k}$ the total number of images seen during the current round, $\eta_{s}$ the server learning rate, $\operatorname{\boldsymbol{w}}^{t}$ the global model and $\operatorname{\boldsymbol{w}}^{t}_{k}$ the local update resulting from training on client $k$’s data with ClientOpt for $E$ epochs. FedAvg [39] computes $\operatorname{\boldsymbol{w}}^{t+1}$ as $\sum_{k\in\operatorname{\mathcal{C}}^{t}}\nicefrac{{N_{k}}}{{N}}\operatorname{\boldsymbol{w}}^{t}_{k}$, corresponding to one SGD step on the pseudo-gradient $\Delta^{t}_{\operatorname{\boldsymbol{w}}}$ with $\eta_{s}=1$ [47].
@@ 3.2 Sharpness-aware Minimization
Sam [12] jointly minimizes the loss value and the sharpness of the loss landscape by solving the min-max problem
@@
$\operatorname*{argmax}_{\|\operatorname{\boldsymbol{\epsilon}}\|\leq\rho}f(\operatorname{\boldsymbol{w}})+\operatorname{\boldsymbol{\epsilon}}^{\top}\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})=\rho\frac{\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})}{\|\nabla_{\operatorname{\boldsymbol{w}}}f(\operatorname{\boldsymbol{w}})\|}\triangleq\operatorname{\hat{\boldsymbol{\epsilon}}}(\operatorname{\boldsymbol{w}}).\vspace{-5pt}$ (5)